# Optimizing Breast Cancer Detection in Mammograms: A Comprehensive Study of Transfer Learning, Resolution Reduction, and Multi-View Classification

### v2

@@ Abstract
Mammography, an X-ray-based imaging technique, plays a crucial role in the early detection of breast cancer. Its accuracy heavily depends on expert radiologists, making it essential to minimize interpretation errors. To support radiologists, various computer-aided detection and diagnostic methods have been proposed, increasingly leveraging advancements in artificial intelligence and machine learning.
Over recent years, mammogram analysis has evolved significantly—from early patch-based classifiers, which examine only localized regions of images, to full-image classifiers, and later towards multi-view systems that simultaneously integrate multiple perspectives of the mammographic exam for enhanced accuracy. Despite this progression, critical questions remain, such as whether multi-view systems consistently outperform single-view approaches.
In this paper, we systematically evaluate and compare the effectiveness of single-view and multi-view mammogram classification techniques. Our research introduces models that achieve superior performance relative to existing state-of-the-art approaches in both single-view and two-view classification scenarios. Furthermore, our findings provide valuable insights into optimal model architectures and effective transfer learning strategies, paving the way for more accurate and efficient mammogram interpretation. The inference code and model are available at https://github.com/dpetrini/multiple-view.
@@ 1 Introduction
Initial research, such as the study by Kooi et al. [3], explored convolutional neural networks (CNNs) for classifying mammograms. However, direct application of CNN architectures designed for natural images did not initially yield promising results due to fundamental differences between natural images and mammograms. Specifically, natural images typically feature prominent objects occupying substantial portions of the image area, whereas cancerous lesions in mammograms are often subtle and localized to small regions. Moreover, natural images contain three color channels (Red, Green, Blue), whereas mammograms are single-channel grayscale images with high resolutions—commonly between 3,000 and 4,000 pixels in both dimensions. These high resolutions significantly increase computational complexity, challenging standard CNN-based classification techniques and conventional transfer learning strategies, thereby motivating the development of specialized methodologies.
@@
Through addressing these questions, we propose enhanced models that achieve state-of-the-art performance for both single-view and two-view mammogram classification tasks. Our findings provide valuable insights into the choice of CNN architectures and transfer learning strategies, ultimately guiding the development of more effective and efficient mammographic analysis systems.
@@ 2 Materials & Methods
For breast cancer classification, machine learning studies may rely on proprietary mammography datasets, as in Wu et al. [7] and McKinney et al. [8], or on publicly accessible datasets. In this work, we exclusively utilize open-access mammographic repositories. In the next topic, we summarize the key characteristics of public datasets available, and in the following topics, we provide a literature review of techniques for breast cancer analysis, covering single view and multi-view architectures. We conclude this section describing the proposed methodology for enhancing mammogram classification performance.
@@ 2.1 Public Datasets
There are two principal types of mammography: screen-film mammography (SFM, conventional film-based) and full-field digital mammography (FFDM). To process SFM digitally, film images must first be scanned—further degrading their quality—whereas FFDM is inherently digital. A simple visual comparison of scanned SFM versus FFDM confirms that the latter offers substantially superior image quality. Accordingly, it is evident that training and evaluating an AI system on native digital mammograms will yield better results than using digitized analogue films.
@@
The CBIS-DDSM [13] is a curated and enhanced subset of the original DDSM, specifically tailored to support research applications. It comprises 3,103 mammographic images, benign or malignant, in this case confirmed by biopsy, and includes detailed lesion annotations to facilitate automated analysis.
@@
VinDr-Mammo [16] is a full-field digital mammography (FFDM) collection of 5,000 four-view exams (20,000 images) with BI-RADS assessments, breast density, and lesion annotations (masses, calcifications, asymmetries, architectural distortions); each exam was independently double-read with arbitration for discordance, and only BI-RADS 3–5 exams include lesion marks.
@@ 2.2 Related work
Below, we outline neural network–based CAD systems for single-view mammogram analysis to address the challenge of mammogram classification.
As mentioned earlier, Shen et al. [4] proposed an strategy based in two transfer-learning stages. They achieved an AUC of 0.87 using an specific selection of CBIS-DDSM images for test set. Other researchers [21] evaluated Shen’s method in the original test set and achieved an AUC of 0.75.
Shu et al. [22] proposed two novel pooling techniques to replace the conventional average-pooling and max-pooling layers, using the same dataset. Their method achieves a maximum AUC of 0.838. However, it is unclear whether they employed the original dataset split, since they report using an 85/15% train/test partition, whereas the original dataset is divided 79/21%.
Wei et al. [21], also using CBIS-DDSM dataset introduced the use of neural-network morphing in lieu of traditional transfer learning. Using the original train/test split, they report AUCs of 0.7964 for the single model without test-time augmentation (TTA), 0.8187 for the single model with TTA, and 0.8313 for an ensemble of four models with TTA. By contrast, they obtain an AUC of 0.9427 when employing a fixed random train/test split; these results are not directly comparable due to different test splits.
@@
Bhat et al. [24] employ a technique called AUCReshaping, which aims to modify the ROC curve to improve evaluation within specific sensitivity and specificity ranges. Using the VinDr-Mammo dataset, they first classify breast density and report a high AUC of 0.97. They also evaluate BI-RADS categories, grouping 1, 2, and 3 as the negative class and 4 and 5 as the positive class, achieving an AUC of 0.77. This grouping yields a different class balance than our further experiments.
@@
Khan et al. [25] introduce a multi-view, patch-based approach using a simplified VGG-style convolutional network to analyze mammographic ROIs from the MIAS and CBIS-DDSM datasets. In three sequential training stages, their model first distinguishes normal from abnormal patches on MIAS with an AUC of 0.934, then classifies CBIS-DDSM patches as masses versus calcifications, AUC of 0.923, and finally differentiates benign from malignant lesions, AUC of 0.769.
@@ 2.3 Methodology
By investigating these questions, we aim to refine best practices in machine learning-based mammogram classification, optimizing both accuracy and computational efficiency. In doing so, we developed models that surpass previous results for both single-view and two-view classifiers on the CBIS-DDSM and VinDr-Mammo dataset. To select the base model architectures, inspired by [30], we will follow an evolutionary scale starting from ResNet [31] until ConvNeXt [32], using open source implementations from Wightman [33] and from PyTorch [34] and will evaluate different training approaches, like patch-based-pre-train, resizing and others, as summarized in Figure 1.
@@
We will use only publicly available mammogram datasets. Our source code and models are available online [35].
@@ 3.2.2 Patch Classifier
The patch classifier is initialized with weights pre-trained on ImageNet and trained using the Adam optimizer, starting with an initial learning rate of $2\times 10^{-5}$. The learning rate follows a “warm-up and cyclic cosine” schedule, configured with a period of 3, a maximum learning rate delta of $2\times 10^{-4}$, and a warm-up phase lasting 4 epochs. The results, summarized in Table 2, demonstrate that among the evaluated base models, ConvNeXt-Base achieves the highest performance.
Table 2: Accuracy of patch classifiers. Models pre-trained on ImageNet-21k or -22k are explicitly marked, while those trained on ImageNet-1k are unmarked.
Model Batch Size Test mean Best test
ViT-Base-CLIP-Laion-32 56 0.6464±0.0109 0.6487
MNASNet-1.0 144 0.6901±0.0269 0.7152
MobileNetV2 192 0.7391±0.0039 0.7369
EfficientNetV2-S (21k) 56 0.7524±0.0056 0.7460
MobileNetV3-Large 96 0.7357±0.0100 0.7462
EfficientNet-B1 192 0.7585±0.0090 0.7495
DenseNet169 144 0.7594±0.0053 0.7540
ResNet-18 192 0.7508±0.0075 0.7579
ResNeXt-50-32x4d 144 0.7580±0.0051 0.7585
DenseNet201 96 0.7617±0.0014 0.7604
EfficientNet-B0 192 0.7585±0.0035 0.7609
EfficientNet-B2 192 0.7550±0.0077 0.7637
EfficientNetV2-M 96 0.7501±0.0129 0.7641
ResNet-101 144 0.7574±0.0087 0.7644
ResNet-50 192 0.7550±0.0087 0.7650
EfficientNet-B4 96 0.7639±0.0057 0.7666
ConvNeXt-Base (22k) 96 0.7740±0.0063 0.7680
DenseNet121 144 0.7685±0.0119 0.7691
EfficientNetV2-S 96 0.7643±0.0111 0.7695
SwinV2-Base-In22k-FT-1k (22k) 56 0.7725±0.0048 0.7695
EfficientNet-B3 144 0.7661±0.0050 0.7700
ConvNeXt-Tiny 96 0.7711±0.0051 0.7721
ConvNeXt-Small 96 0.7791±0.0081 0.7784
ConvNeXt-Base 96 0.7831±0.0082 0.7918
@@ 3.2.3 Patch Based Classifier (PBC)
The resulting AUCs are presented in Table 3. Notably, we observed that the best-performing base models for patch classification did not consistently perform well for full mammogram classification. Specifically, while ConvNeXt-Base was the top-performing model for patch classification, EfficientNet-B3 emerged as the best base model for classifying whole single-view mammograms. In this table and in following ones, “Train size” indicates the resolution of the pre-training images of the base model.
Table 3: AUCs obtained by single-view patch-based classifiers (PBC). PBC/DC indicates which approach generates the highest AUC in the best test, comparing with Table 4.
Model Train size Test mean Best test PBC /DC
[…]
@@ 3.2.4 Direct Classifier (DC)
We trained the single-view classifiers using transfer learning directly from base models pre-trained on ImageNet, bypassing the patch classifier. The resulting AUCs are summarized in Table 4. Consistent with previous findings, EfficientNet-B3 achieved the best performance among the evaluated base models.
Table 4: AUCs obtained by single-view direct classifiers (DC). PBC/DC indicates which approach generates the highest AUC in the best test, comparing with Table 3.
Model Train size Test mean Best test PBC /DC
DenseNet169 224 0.6784±0.0184 0.6524±0.0222 PBC
DenseNet121 224 0.6697±0.0057 0.6696±0.0219 PBC
ConvNeXt-Base 224 0.6817±0.0178 0.6701±0.0219 PBC
ConvNeXt-Tiny 224 0.6889±0.0123 0.6791±0.0218 PBC
ConvNeXt-Small 224 0.6674±0.0152 0.6802±0.0217 PBC
DenseNet201 224 0.6878±0.0078 0.6812±0.0217 PBC
ResNeXt-50-32x4d 224 0.6876±0.0054 0.6909±0.0215 PBC
ResNet-18 224 0.6883±0.0073 0.6935±0.0215 PBC
ConvNeXt-Base (22k) 224 0.7019±0.0262 0.7289±0.0207 PBC
MNASNet-1.0 224 0.7633±0.0147 0.7427±0.0203 PBC
MobileNetV2 224 0.7584±0.0251 0.7659±0.0196 DC
ResNet-101 224 0.7698±0.0092 0.7702±0.0195 PBC
EfficientNet-B4 384 0.8023±0.0064 0.7933±0.0187 DC
ResNet-50 224 0.7293±0.0598 0.7993±0.0185 DC
EfficientNetV2-M 384 0.7605±0.0376 0.8024±0.0183 PBC
EfficientNet-B2 260 0.8058±0.0071 0.8078±0.0181 PBC
EfficientNet-B1 240 0.8023±0.0156 0.8117±0.0180 DC
EfficientNet-B0 224 0.7691±0.0525 0.8127±0.0179 DC
MobileNetV3-Large 224 0.7792±0.0247 0.8132±0.0179 DC
EfficientNetV2-S 300 0.8100±0.0251 0.8266±0.0174 DC
EfficientNetV2-S (21k) 300 0.8098±0.0162 0.8301±0.0172 DC
EfficientNet-B3 300 0.8087±0.0177 0.8313±0.0172 PBC
@@ 3.2.5 Conclusions about Patch Classifier
Comparing Tables 3 and 4, we observe that PBC and DC exhibit similar performance. Both approaches achieved their best results using the EfficientNet-B3 network, with AUC values of 0.8325±0.0171 for PBC and 0.8313±0.0172 for DC — both very close (Table 5).
@@ 3.2.6 Conclusions about Base-Models
In this section, we analyze Tables 3 and 4 to address the question (2), concluding:
@@
  • Larger training image sizes (300×300 or greater) were associated with the best-performing networks for both PBC (Table 3) and DC (Table 4).
@@ 3.2.7 Comparison with Previous Work
3.3 Classifiers with Resized in CBIS-DDSM
@@ 3.3.1 Fixed Resizing Classifier (FRC)
We downscaled the original 1152×892 images to 576×448 using the INTER_AREA interpolation method from the OpenCV library. Transfer learning was then performed directly from the ImageNet pre-trained weights. The results, presented in Table 6, show that the highest AUC achieved was 0.8167±0.0178. This performance is notably lower than the results obtained without downsampling: 0.8313±0.0172 for direct classification (DC).
Table 6: AUCs of single-view direct classifiers using fixed resizing (FRC).
Model Train size Test mean Best test
[…]
@@ 3.3.2 Learn-to-Resize Classifier (LRC)
Talebi and Milanfar [29] proposed a technique called “Learn to Resize,” which integrates bilinear resizing with convolutional layers, enabling the model to optimize resizing for improved classification performance. As shown in Table 7, the highest AUC achieved using this method was 0.7958±0.0186, which is lower than the results obtained with fixed resizing (FRC).
Table 7: AUCs of single-view direct classifiers with learned resizing (LRC).
Model Train size Test mean Best test
[…]
@@ 3.3.3 Conclusions on Resizing
From Tables 6 and 7, we observed that the best-performing networks using FRC and LRC utilized base models pre-trained on small image sizes of 224×224.
@@ 3.6.1 Patch Classifier and Base-Model
To answer questions (1) and (2) for VinDr-Mammo dataset, we prepared the patches similar to Section 3.2.2, considering four categories: background and the lesions with Bi-Rads 3, 4 and 5. The backbone model with the best performance was ConvNeXt_Base_22k (Table 13).

### v3

@@ Abstract
Mammography, an X-ray–based imaging technique, remains central to the early detection of breast cancer. Recent advances in artificial intelligence have enabled increasingly sophisticated computer-aided diagnostic methods, evolving from patch-based classifiers to whole-image approaches and then to multi-view architectures that jointly analyze complementary projections. Despite this progress, several critical questions remain unanswered.
In this study, we systematically investigate these issues by addressing five key research questions: (1) the role of patch classifiers in performance, (2) the transferability of natural-image-trained backbones, (3) the advantages of “learn-to-resize” over conventional downscaling, (4) the contribution of multi-view integration, and (5) the robustness of findings across varying image quality. Beyond benchmarking, our experiments demonstrate clear performance gains over prior work. For the CBIS-DDSM dataset, we improved single-view AUC from 0.8153 to 0.8343, and multiple-view AUC from 0.8483 to 0.8658. Using a new comparative method, we also observed a 0.0217 AUC increase when extending from single- to multiple-view analysis. On the complete VinDr-Mammo dataset, the multiple-view approach further improved results, achieving a 0.0492 AUC increase over single view and reaching 0.8511 AUC overall.
These results establish new state-of-the-art benchmarks, providing clear evidence of the advantages of multi-view architectures for mammogram interpretation. Beyond performance, our analysis offers principled insights into model design and transfer learning strategies, contributing to the development of more accurate and reliable breast cancer screening tools. The inference code and trained models are publicly available at https://github.com/dpetrini/multiple-view.
Keywords – deep learning, mammography, breast cancer, CBIS-DDSM, VinDr-Mammo, multiple-view classifier.
@@ 1 Introduction
Early applications of convolutional neural networks (CNNs) to mammogram classification, such as the work of Kooi et al. [3], highlighted the challenges of direct transfer from natural image tasks. The subtle and localized appearance of cancerous lesions, combined with the high-resolution, single-channel nature of mammograms, limits the effectiveness of conventional CNNs and underscores the need for specialized methodologies.
@@
Through addressing these questions, we propose enhanced models that achieve state-of-the-art performance for both single-view and multiple-view mammogram classification tasks. Our findings provide valuable insights into the choice of CNN architectures and transfer learning strategies, ultimately guiding the development of more effective and efficient mammographic analysis systems.
@@ 2 Materials & Methods
For breast cancer classification, machine learning studies may rely on proprietary mammography datasets, as in Wu et al. [7] and McKinney et al. [8], or on publicly accessible datasets. In this work, we exclusively utilize open‐access mammographic repositories. In the next topic, we summarize the key characteristics of public datasets available, and in the following topics, we provide a literature review of techniques for breast cancer analysis, covering single view and multi-view architectures. We conclude this section describing the proposed methodology for enhancing mammogram classification performance.
@@ 2.1 Public Datasets
There are two principal types of mammography: screen-film mammography (SFM, conventional film-based) and full-field digital mammography (FFDM). To process SFM digitally, film images must first be scanned—further degrading their quality—whereas FFDM is inherently digital. A simple visual comparison of scanned SFM versus FFDM confirms that the latter offers substantially superior image quality, as in Figure 1. Accordingly, it is evident that training and evaluating an AI system on native digital mammograms will yield better results than using digitized analogue films.
(a) Mammogram in the left, and mask in the right. The mask indicates the location of the lesion. In rightmost bottom, the patch of mass lesion.
(b) Mammogram with a box (from metadata) indicating the location of lesion. In leftmost bottom, the patch of mass lesion.
Figure 1: Samples of public dataset used in this work. CBIS-DDSM film-based capture on the left and VinDr-Mammo digital capture on the right.
@@
The CBIS‐DDSM [13] is a curated and enhanced subset of the original DDSM, specifically tailored to support research applications. It comprises 3,103 mammographic images, benign or malignant, in this case confirmed by biopsy, and includes detailed lesion annotations to facilitate automated analysis.
@@
VinDr-Mammo [16] is a full-field digital mammography (FFDM) collection of 5,000 four-view exams (20,000 images) with BI-RADS assessments, breast density, and lesion annotations (masses, calcifications, asymmetries, etc); each exam was independently double-read with arbitration for discordance, and only BI-RADS 3–5 exams include lesion marks.
@@ 2.2 Related work
Below, we outline neural network–based CAD systems for single‐view mammogram analysis to address the challenge of mammogram classification.
As mentioned earlier, Shen et al. [4] proposed an strategy based in two transfer‐learning stages. They achieved an AUC of 0.87 using an specific selection of CBIS-DDSM images for test set. Other researchers [21] evaluated Shen’s method in the original test set and achieved an AUC of 0.75.
Shu et al. [22] proposed two novel pooling techniques to replace the conventional average‐pooling and max‐pooling layers, using the same dataset. Their method achieves a maximum AUC of 0.838. However, it is unclear whether they employed the original dataset split, since they report using an 85/15% train/test partition, whereas the original dataset is divided 79/21%.
Wei et al. [21], also using CBIS-DDSM dataset, introduced the use of neural-network morphing in lieu of traditional transfer learning. Using the original train/test split, they report AUCs of 0.7964 for the single model without test-time augmentation (TTA), 0.8187 for the single model with TTA, and 0.8313 for an ensemble of four models with TTA. By contrast, they obtain an AUC of 0.9427 when employing a fixed random train/test split; these results are not directly comparable due to different test splits.
@@
Bhat et al. [24] employ a technique called AUCReshaping, which aims to modify the ROC curve to improve evaluation within specific sensitivity and specificity ranges. Using the VinDr‐Mammo dataset, they evaluated two BI-RADS-based categories, grouping 1, 2, and 3 as the negative class and 4 and 5 as the positive class, achieving an AUC of 0.77. This grouping yields a different class balance than our further experiments.
Shah et al. [25] explored backbones architectures as AlexNet [26], ResNet[27], DenseNet [28] and EfficientNet [29] using a dataset composed of DDSM [9], RSNA [18] and synthetic images, totaling 6210 mammograms. Their experiments reported an AUC of 94% in both EfficientNet and DenseNet.
@@
Khan et al. [30] introduce a multi‐view, patch‐based approach using a simplified VGG‐style convolutional network to analyze mammographic ROIs from the MIAS and CBIS‐DDSM datasets. In three sequential training stages, their model first distinguishes normal from abnormal patches on MIAS with an AUC of 0.934, then classifies CBIS‐DDSM patches as masses versus calcifications, AUC of 0.923, and finally differentiates benign from malignant lesions, AUC of 0.769.
@@
Shah et al. [34] proposed a two-view architecture in which mammographic images were processed through EfficientNet backbones, followed by feature fusion. Using the RSNA dataset, they achieved an accuracy of 0.99, outperforming other comparative methods.
@@ 2.3 Methodology
By investigating these questions, we aim to refine best practices in machine learning-based mammogram classification, optimizing both accuracy and computational efficiency. In doing so, we developed models that surpass previous results for both single-view and two-view classifiers on the CBIS-DDSM and VinDr-Mammo public datasets. To select the base model architectures, inspired by [36], we will follow an evolutionary scale starting from ResNet [27] until ConvNeXt [37], using open source implementations from Wightman [38] and from PyTorch [39] and will evaluate different training approaches, like patch-based-pre-train, resizing and others, as summarized in Figure 2.
@@ 3.2.2 Patch Classifier
The patch classifier is initialized with weights pre-trained on ImageNet and trained using the Adam optimizer, starting with an initial learning rate of $2\times 10^{-5}$. The learning rate follows a “warm-up and cyclic cosine” schedule, configured with a period of 3, a maximum learning rate delta of $2\times 10^{-4}$, and a warm-up phase lasting 4 epochs. The results, summarized in Table S1 (Data Supplement S1), demonstrate that among the evaluated base models, ConvNeXt-Base achieves the highest performance, with an accuracy of 0.7918.
@@ 3.2.3 Patch Based Classifier (PBC)
The resulting AUCs are presented in Table S2. Notably, we observed that the best-performing base models for patch classification did not consistently perform well for full mammogram classification. Specifically, while ConvNeXt-Base was the top-performing model for patch classification, EfficientNet-B3, with and AUC of 0.8325±0.0171, emerged as the best base model for classifying whole single-view mammograms.
@@ 3.2.4 Direct Classifier (DC)
We trained the single-view classifiers using transfer learning directly from base models pre-trained on ImageNet, bypassing the patch classifier. The resulting AUCs are summarized in Table S3. Consistent with previous findings, EfficientNet-B3 achieved the best performance among the evaluated base models, with an AUC of 0.8313±0.0172.
@@ 3.2.5 Conclusions about Patch Classifier
As shown in Table 2, PBC and DC yield comparable performance when implemented with the EfficientNet-B3 backbone.
@@ 3.2.6 Conclusions about Base-Models
In this section, we analyze Tables S2 and S3 to address the question (2), concluding:
@@
  • Larger training image sizes (300×300 or greater) were associated with the best-performing networks for both PBC (Table S2) and DC (Table S3).
@@ 3.2.7 Comparison with Previous Work
3.3 Classifiers with Resized images in CBIS-DDSM
@@ 3.3.1 Fixed Resizing Classifier (FRC)
We downscaled the original 1152×892 images to 576×448 using the INTER_AREA interpolation method from the OpenCV library. Transfer learning was then performed directly from the ImageNet pre-trained weights. The results, presented in Table S4, show that the highest AUC achieved was 0.8167±0.0178. This performance is notably lower than the results obtained without downsampling: 0.8313±0.0172 for direct classification (DC).
@@ 3.3.2 Learn-to-Resize Classifier (LRC)
Talebi and Milanfar [35] proposed a technique called “Learn to Resize,” which integrates bilinear resizing with convolutional layers, enabling the model to optimize resizing for improved classification performance. As shown in Table S5, the highest AUC achieved using this method was 0.7958±0.0186, which is lower than the results obtained with fixed resizing (FRC).
@@ 3.3.3 Conclusions on Resizing
In our tests, larger images yielded better results; however, a different study [44] demonstrated that higher resolutions (2,304 × 1,792) did not improve performance.
From Tables S4 and S5, we observed that the best-performing networks using FRC and LRC utilized base models pre-trained on small image sizes of 224×224.
@@ 3.6.1 Patch Classifier and Base-Model
To answer questions (1) and (2) for VinDr-Mammo dataset, we prepared the patches similar to Section 3.2.2, considering four categories: background and the lesions with Bi-Rads 3, 4 and 5. The backbone model with the best performance was ConvNeXt_Base_22k with an accuracy of 0.7052 (Table S6).
3.6.2 Patch Based Classifier (PBC)
Using the weights obtained from the patch classifier, we constructed single-view patch-based classification (PBC) models and achieved the results presented in Table S7. Training the the best model, ConvNeXt-Base (22k) that achieved an AUC of 0.8510±0.0163, required approximately 24 hours for the three rounds on an NVIDIA A100 GPU with 40 GB of memory, making it the most time-intensive single-view model in this study.
3.6.3 Direct Classifier (DC)
The results of the Direct Classifier (DC) experiments are summarized in Table S8. DenseNet169 emerged as the top-performing base model, with an AUC of 0.8134±0.0136. Notably, the best-performing Patch-Based Classifier (PBC), ConvNeXt_Base_22k, showed a significant drop in performance when applied to the DC approach. Conversely, DenseNet169, which excelled in the DC task, performed poorly in the PBC setting
3.6.4 Conclusions about PBC
We compared both results shown above by running the DeLong test and obtained the value $p$ = 0.0013. This indicates with statistical significance that the best PBC model (ConvNeXt-Base (22k)) is superior to the best DC model (DenseNet16) for classifying high-quality mammograms. Due to this result, we recommend the PBC approach for the VinDr-Mammo set with 100% digital mammograms, despite the fact that it requires more training time.
On the other hand, four base models performed better in the Patch-Based Classifier (PBC) approach, while three models showed superior performance in the Direct Classifier (DC) approach (see Fig. 5 and Tables S7 and S8). Based on these results, it is not possible to conclude that the PBC approach is universally superior to the DC approach for classifying high-quality mammograms, as performance varies depending on the base model.
3.6.5 Conclusions on Base Models
3.7 Classifiers with Resized Images in VinDr-Mammo
To answer the question (3) for VinDr-Mammo, in the same way as the CBIS-DDSM, we downsample mammograms to half their size before classifying them.
Consistent with Section 3.3.3, the best Learn-to-Resize classifier (LRC; full results in Table S9) underperformed the best Fixed-Resizing classifier (FRC; full results in Table S10). Since LRC is both computationally more complex and resource-intensive than FRC, the Learn-to-Resize approach offers no practical benefit.
3.8 Two-View Classifiers in VinDr-Mammo
To address question (4), we followed the same network creation and training process outlined in previous sections. We constructed a two-view classifier using the ConvNeXt_Base_22k network, initializing it with weights from the single-view PBC approach. Additionally, we processed each view (CC and MLO) individually using the best-performing classifier from the previous section (ConvNeXt_Base_22k) and combined the results using both average and maximum operations. The results are shown in Table 8.
Model Two-View Average Maximum
[…]
4 Conclusions
[…]
Supporting Information
Additionalsupporting information can be found online in the Supporting Information section. Data Supplement S1. Supplemental material.
Data Supplement S1. Supplemental material.
Table S1. Patch Classifier results
Table S1: Accuracy of patch classifiers. Models pre-trained on ImageNet-21k or -22k are explicitly marked, while those trained on ImageNet-1k are unmarked.
Model Batch Size Test mean Best test
ViT-Base-CLIP-Laion-32 56 0.6464±0.0109 0.6487
MNASNet-1.0 144 0.6901±0.0269 0.7152
MobileNetV2 192 0.7391±0.0039 0.7369
EfficientNetV2-S (21k) 56 0.7524±0.0056 0.7460
MobileNetV3-Large 96 0.7357±0.0100 0.7462
EfficientNet-B1 192 0.7585±0.0090 0.7495
DenseNet169 144 0.7594±0.0053 0.7540
ResNet-18 192 0.7508±0.0075 0.7579
ResNeXt-50-32x4d 144 0.7580±0.0051 0.7585
DenseNet201 96 0.7617±0.0014 0.7604
EfficientNet-B0 192 0.7585±0.0035 0.7609
EfficientNet-B2 192 0.7550±0.0077 0.7637
EfficientNetV2-M 96 0.7501±0.0129 0.7641
ResNet-101 144 0.7574±0.0087 0.7644
ResNet-50 192 0.7550±0.0087 0.7650
EfficientNet-B4 96 0.7639±0.0057 0.7666
ConvNeXt-Base (22k) 96 0.7740±0.0063 0.7680
DenseNet121 144 0.7685±0.0119 0.7691
EfficientNetV2-S 96 0.7643±0.0111 0.7695
SwinV2-Base-In22k-FT-1k (22k) 56 0.7725±0.0048 0.7695
EfficientNet-B3 144 0.7661±0.0050 0.7700
ConvNeXt-Tiny 96 0.7711±0.0051 0.7721
ConvNeXt-Small 96 0.7791±0.0081 0.7784
ConvNeXt-Base 96 0.7831±0.0082 0.7918
Table S2. Patch Based Classifier (PBC) results
Table S2: AUCs obtained by single-view patch-based classifiers (PBC). PBC/DC indicates which approach generates the highest AUC in the best test, comparing with Table S3. In this table and in following ones, “Train size” indicates the resolution of the pre-training images of the base model.
Model Train size Test mean Best test PBC /DC
[…]
Table S3. Direct Classifier (DC) results
Table S3: AUCs obtained by single-view direct classifiers (DC). PBC/DC indicates which approach generates the highest AUC in the best test, comparing with Table S2.
Model Train size Test mean Best test PBC /DC
DenseNet169 224 0.6784±0.0184 0.6524±0.0222 PBC
DenseNet121 224 0.6697±0.0057 0.6696±0.0219 PBC
ConvNeXt-Base 224 0.6817±0.0178 0.6701±0.0219 PBC
ConvNeXt-Tiny 224 0.6889±0.0123 0.6791±0.0218 PBC
ConvNeXt-Small 224 0.6674±0.0152 0.6802±0.0217 PBC
DenseNet201 224 0.6878±0.0078 0.6812±0.0217 PBC
ResNeXt-50-32x4d 224 0.6876±0.0054 0.6909±0.0215 PBC
ResNet-18 224 0.6883±0.0073 0.6935±0.0215 PBC
ConvNeXt-Base (22k) 224 0.7019±0.0262 0.7289±0.0207 PBC
MNASNet-1.0 224 0.7633±0.0147 0.7427±0.0203 PBC
MobileNetV2 224 0.7584±0.0251 0.7659±0.0196 DC
ResNet-101 224 0.7698±0.0092 0.7702±0.0195 PBC
EfficientNet-B4 384 0.8023±0.0064 0.7933±0.0187 DC
ResNet-50 224 0.7293±0.0598 0.7993±0.0185 DC
EfficientNetV2-M 384 0.7605±0.0376 0.8024±0.0183 PBC
EfficientNet-B2 260 0.8058±0.0071 0.8078±0.0181 PBC
EfficientNet-B1 240 0.8023±0.0156 0.8117±0.0180 DC
EfficientNet-B0 224 0.7691±0.0525 0.8127±0.0179 DC
MobileNetV3-Large 224 0.7792±0.0247 0.8132±0.0179 DC
EfficientNetV2-S 300 0.8100±0.0251 0.8266±0.0174 DC
EfficientNetV2-S (21k) 300 0.8098±0.0162 0.8301±0.0172 DC
EfficientNet-B3 300 0.8087±0.0177 0.8313±0.0172 PBC
Table S4. Fixed Resizing Classifier (FRC) results
Table S4: AUCs of single-view direct classifiers using fixed resizing (FRC).
Model Train size Test mean Best test
[…]
Table S5. Learn-to-Resize Classifier (LRC) results
Table S5: AUCs of single-view direct classifiers with learned resizing (LRC).
Model Train size Test mean Best test
[…]
Table S6. Patch Classifier and Base-Model (VinDr-Mammo) results